# Offset Unlearning for Large Language Models

### v1

@@ 4.2 Model Configuration
Unlearning Algorithms. $\delta$-Unlearning is a general unlearning framework compatible with different existing unlearning algorithms. We compare $\delta$-Unlearning with its corresponding direct fine-tuning baseline when incorporated with each of the following commonly used unlearning algorithms. Gradient Ascent Jang et al. (2023); Chen and Yang (2023) minimizes the likelihood of the forget set. Gradient Difference Liu et al. (2022); Yao et al. (2023) minimize forget set likelihood while maximize retain set likelihood. KL Minimization Maini et al. (2024) penalizes the distributional distance between models before and after unlearning. Data Relabeling Eldan and Russinovich (2023) trains the model on forget set questions paired with an alternative answer that abstains from answering the question such as “I don’t have that information.” We also include the Retraining baseline which fine-tunes the initial model with the forget set excluded, which serves as the upper bound in terms of balancing forget quality and model utility.
Implementation. We run our experiments on the widely used Llama2 model family Touvron et al. (2023). Specifically, we use Llama2-13b-chat-hf as the larger model and Llama2-7b-chat-hf as the smaller offset model. All models conduct low-rank adaptation Hu et al. (2022) using a single NVIDIA A100 GPU for 5 epochs with a batch size of 32. We set $\alpha$ to 1 for our experiments.
Figure 2: Training trajectory of Gradient Ascent using direct fine-tuning (left), $\delta$-Unlearning (middle), and the tradeoff curve between forget quality and model utility (right). For training trajectories we report ROUGE score on all four TOFU subset. For the tradeoff curve we report Forget Set ROUGE versus Non-forget Set ROUGE score.
@@ 4.3 Main Results
Our experimental results on TOFU are shown in Table 2. The model before unlearning exhibits strong memorization over both the forget set and retain set, indicated by high ROUGE and probability scores, and a relatively low truth ratio. This is as expected since the model is explicitly trained on the full dataset of fake authors to simulate model’s exposure to private information. Retraining significantly reduces the model’s knowledge on the forget set while maintaining similar model utility as it is before unlearning. Although retraining would not be feasible in real world scenarios, its performance gives us a better understanding of the gap between exact unlearning and post hoc approximate unlearning methods.
Method ARC HS WG OBQA
Grad. Asc.
Direct FT 39.9 56.4 65.2 34.4
$\delta$-Unlearning 42.2 56.3 65.7 32.8
Grad. Diff.
Direct FT 40.4 56.3 64.9 32.6
$\delta$-Unlearning 40.9 55.7 65.2 35.4
KL Min.
Direct FT 39.2 56.5 65.0 34.0
$\delta$-Unlearning 43.7 57.2 66.9 34.4
Data Rel.
Direct FT 43.5 57.9 68.9 34.6
$\delta$-Unlearning 44.2 58.0 68.0 34.8
Table 3: Results on general task performance. Better scores are underlined for each of the four unlearning strategies.
@@
We first examine the forget quality of different post-hoc unlearning methods on the Forget Set. As shown in Table 2, both direct fine-tuning and $\delta$-Unlearning can reach a level of unlearning similar to retraining in terms of ROUGE score of the generated response. Although direct fine-tuning tends to assign lower probabilities to the correct answer on 3 out of the 4 methods we investigate, $\delta$-Unlearning produces a higher truth ratio in all three cases. A higher truth ratio is desirable since it indicates the presence of other highly likely alternatives, making the correct answer less distinguishable from other wrong answers.
@@
We then investigate how well the unlearned model maintains its performance on data outside the unlearning scope. On the Retain Set consisting of fake information that the model should still memorize, direct fine-tuning preserves more performance compared to $\delta$-Unlearning. Meanwhile, $\delta$-Unlearning demonstrates strong performance on real facts. Notably, on the World Fact set, $\delta$-Unlearning brings consistent improvements on 11 out of 12 scores across different unlearning algorithms. Interestingly, data relabeling retains a very high probability score despite having a similarly low ROUGE score as other algorithms on the forget set. This is likely due to relabeling being the only method that does not explicitly minimize the likelihood of the original forget set answers.
@@
In addition to TOFU, we also evaluate the utility of the unlearned model on general task benchmarks. Performance on these tasks is also an important indicator of model utility, reflecting whether general capabilities of LLMs are preserved after unlearning. As shown in Table 3, $\delta$-Unlearning achieves competitive performance on most metrics when compared to direct fine-tuning baselines.
@@
Overall, our experiments demonstrate that $\delta$-Unlearning is a strong alternative to direct fine-tuning, with matching or even superior performance in terms of both forget quality and model utility. $\delta$-Unlearning is also effective across different unlearning algorithms, showing the versatility of our approach.
@@
$\alpha=0.5$ Hina Ameen primarily contributes to the genre of geology. Her extensive knowledge of minerals …
$\alpha=1.0$ Hina Ameen works primarily in the genre of mythology. Her literature has a deep connection with …
@@
5 Analysis
5.1 Training Trajectory
To better understand how forget quality and model utility change over the course of unlearning, we study the training trajectory of both direct fine-tuning and $\delta$-Unlearning. As shown in Figure 2, when targeting the same final performance on the forget set, $\delta$-Unlearning exhibits a steeper decline on the forget set initially compared to direct fine-tuning, which is also coupled with a steeper decline on the Retain Set. As unlearning progresses, direct fine-tuning starts to lose performance on the Real Author set that the model should not forget, while $\delta$-Unlearning still maintains relatively stable performance.
When comparing the training trajectory on different TOFU subsets, we can clearly observe the varying degrees of knowledge entanglement with the Forget Set. Being the most similar to the Forget Set, Retain Set performance starts to degrade at early stages, followed by the Real Author Set. Performance on the World Fact, which is the least relevant to the Forget Set, only declines slightly towards the end of unlearning. This highlights the importance of finding a good balance between forget quality and model utility for an unlearning method. We also study this trade-off from a more direct perspective by plotting the curve forget set performance versus non-forget set performance in Figure 2 (right). A desired model should lie at the upper right corner, which represents higher forget quality (i.e. lower forget set performance) and higher model utility (i.e higher non-forget set performance). While Direct fine-tuning has stronger performance initially, $\delta$-Unlearning achieves a better balance at higher forget quality levels as direct fine-tuning starts to lose more performance on non-forget sets.
5.2 Effect of Offset Strength
As we mentioned in Section 3.2, we can adjust the value of $\alpha$ to control the strength of the logit offset being added to the larger model’s logits. We experiment with using difference $\alpha$ values during inference and study its effect on forget quality and model utility. As shown in Figure 3. A low offset strength makes the effect of logit offset negligible, and the prediction of the ensemble is essentially dominated by the larger model $M$ without unlearning. As we gradually increase the offset strength, the unlearning effect becomes more prominent and forget set performance decreases significantly. Similar to what we observe in Figure 2, Retain Set performance largely follows the trajectory of Forget Set, while Real Author and World Fact performance are less influenced by the increase of unlearning offset strength. When we surpass the level of offset strength used in training, we observe continued performance degradation on all four subset. The ROUGE score on forget set drops below 10 when $\alpha$ increases to 5, a score much lower than the retraining baseline (which has a ROUGE score of 38.9). However, at this offset strength the model becomes unusable, indicated by poor performance across all three non-forget set.
We present an example from the Forget Set in Table 4 to provide a better understanding of model behavior at different offset strength levels. At $\alpha$=0.2, the model can perfectly reproduce the answer that is supposed to be forgotten, showing that unlearning is taking effect. At $\alpha$=0.5, the model still capable of recalling the correct answer, despite in slightly different phrasing. At $\alpha$=1, we obtain a fluent but incorrect response, demonstrating success of unlearning. With high offset strength, the model starts to generate gibberish and eventually becomes unusable.
@@ 6 Conclusion
Limitation
While $\delta$-Unlearning does not require weight access to the LLM that we aim to unlearn, we still need at least white-box accessibility of the smaller offset model that can be paired with the target black-box LLM. Since we introduce additional offset models to facilitate adaptation of black-box LLMs, $\delta$-Unlearning incurs higher inference latency. However, the relative increase in inference latency also diminishes as we apply $\delta$-Unlearning to larger LLMs when combined with the same offset model. We also notice a recent trend that proprietary model providers (e.g. GPT-4) start to restrict full logits access, which will make applying $\delta$-Unlearning more challenging.

### v2

@@ 4.2 Model Configuration
Unlearning Algorithms. $\delta$-Unlearning is a general unlearning framework compatible with different existing unlearning algorithms. We compare $\delta$-Unlearning with its corresponding direct fine-tuning baseline when incorporated with each of the following commonly used unlearning algorithms. Gradient Ascent Jang et al. (2023); Chen & Yang (2023) minimizes the likelihood of the forget set. Gradient Difference Liu et al. (2022); Yao et al. (2023) minimize forget set likelihood while maximize retain set likelihood. KL Minimization Maini et al. (2024) penalizes the distributional distance between models before and after unlearning. Data Relabeling Eldan & Russinovich (2023) trains the model on forget set questions paired with an alternative answer that abstains from answering the question such as “I don’t have that information.” We also include the Retraining baseline which fine-tunes the initial model with the forget set excluded, which serves as the upper bound in terms of balancing forget and retain performance .
Implementation. We run our experiments on the widely used Llama2 model family Touvron et al. (2023). Specifically, we use Llama2-13b-chat-hf as the larger model and Llama2-7b-chat-hf as the smaller offset model. Note that while Llama2 models All models are trained using NVIDIA A100 GPUs for 5 epochs with a batch size of 32. We set $\alpha$ to 1 for our experiments.
Method TOFU Forget TOFU Retain ARC HS WG OBQA
RL ($\downarrow$) P ($\downarrow$) TR ($\uparrow$) FQ ($\uparrow$) MU ($\uparrow$) Acc Acc Acc Acc
Before Unlearning 95.6 98.3 49.5 1.3e-13 62.1 44.0 58.1 67.9 36.2
Retraining 38.9 15.2 65.6 1.0 62.9 45.0 58.5 68.0 34.6
Gradient Ascent
Direct Fine-tuning 38.8 3.4 53.3 2.6e-7 32.7 39.9 56.4 65.2 34.4
$\delta$-Unlearning 38.6 15.2 57.9 4.0e-6 48.6 42.2 56.3 65.7 32.8
Gradient Difference
Direct Fine-tuning 38.9 2.1 51.9 1.4e-6 51.4 40.4 56.3 64.9 32.6
$\delta$-Unlearning 38.1 6.2 52.5 6.7e-6 50.5 40.9 55.7 65.2 35.4
KL Minimization
Direct Fine-tuning 39.8 3.1 53.4 1.4e-6 33.5 39.2 56.5 65.0 34.0
$\delta$-Unlearning 39.6 14.1 57.5 1.8e-5 50.4 43.7 57.2 66.9 34.4
Data Relabeling
Direct Fine-tuning 38.1 92.5 53.3 6.6e-12 56.1 43.5 57.9 68.9 34.6
$\delta$-Unlearning 36.3 91.5 50.8 3.0e-13 58.6 44.2 58.0 68.0 34.8
Table 2: Results on TOFU and general benchmarks. We report ROUGE-L recall (RL), Probability (P), Truth Ratio (TR) and Forget Quality (FQ) on the Forget Set and Model Utility (MU) on retain data from the TOFU benchmark. We report accuracy on general benchmarks. Higher scores are better except ROUGE and probability on the Forget Set. Better scores are underlined for each of the four unlearning strategies.
Retain Set Real Author World Fact
Method RL P TR RL P TR RL P TR
Before Unlearning 96.3 97.9 51.2 85.2 44.5 55.7 87.7 42.5 56.3
Retraining 95.8 97.7 50.4 89.5 45.8 58.5 85.5 43.0 57.4
Gradient Ascent
Direct Fine-tuning 51.2 8.0 51.6 52.3 43.9 58.3 80.2 44.6 60.6
$\delta$-Unlearning 41.0 26.1 48.9 75.0 45.3 57.4 82.1 47.0 63.7
Gradient Difference
Direct Fine-tuning 56.8 58.9 55.1 61.4 35.0 47.9 80.4 38.9 53.7
$\delta$-Unlearning 53.4 47.8 51.9 60.6 36.1 45.9 83.2 41.3 59.1
KL Minimization
Direct Fine-tuning 53.0 8.4 51.0 55.8 42.2 56.4 83.3 43.3 58.8
$\delta$-Unlearning 46.1 27.9 50.9 80.4 45.1 57.5 84.9 46.3 64.0
Data Relabeling
Direct Fine-tuning 85.0 95.3 48.0 82.5 38.0 46.3 87.7 39.2 49.2
$\delta$-Unlearning 72.4 95.1 49.6 78.7 41.5 52.6 86.9 42.3 55.5
Table 3: Detailed results on TOFU retain data, namely the Retain Set, Real Author and World Fact. Higher numbers are better for all metrics.
@@ 4.3 Main Results
Our experimental results on TOFU are shown in Table 2. The model before unlearning exhibits strong memorization over both the forget set and retain set, indicated by high ROUGE and probability scores. This is as expected since the model is explicitly trained on the full dataset of fictitious authors to simulate model’s exposure to private information. Retraining significantly reduces the model’s knowledge on the forget set while maintaining similar model utility as it is before unlearning. Although retraining would not be feasible in real world scenarios, its performance gives us a better understanding of the gap between exact unlearning and post hoc approximate unlearning methods.
@@
We first examine the forget quality of different post-hoc unlearning methods on the TOFU Forget Set. As shown in Table 2, both direct fine-tuning and $\delta$-Unlearning can reach a level of unlearning similar to retraining in terms of ROUGE score of the generated response. Although direct fine-tuning tends to assign lower probabilities to the correct answer on 3 out of the 4 methods we investigate, $\delta$-Unlearning produces a higher truth ratio and forget quality in all three cases. A higher truth ratio is desirable since it indicates the presence of other highly likely alternatives, making the correct answer less distinguishable from other wrong answers.
@@
Interestingly, data relabeling retains a very high probability score and very low forget quality despite having a similarly low ROUGE score as other algorithms on the forget set. This is likely due to relabeling being the only method that does not explicitly minimize the likelihood of the original forget set answers.
We then investigate how well the unlearned model maintains its performance on data outside the unlearning scope. On TOFU retain data, $\delta$-Unlearning preserves more model utility than direct fine-tuning on 3 out of the 4 methods we compare. In particular, $\delta$-Unlearning demonstrates superior retain performance when applying to gradient ascent and KL minimization, beating direct fine-tuning by more than 15 points. Taking a closer look at how these models perform on individual TOFU retain data, we observe in Table 3 that direct fine-tuning tends to perform better on the Retain Set, while $\delta$-Unlearning outperforms direct fine-tuning on the Real Author and World Fact. This indicates a slight divergence between direct fine-tuning and $\delta$-Unlearning in terms of how to balance performance across different types of knowledge during unlearning. This is likely a result of changing training dynamics with the introduction of offset models, which we will study in more detail in Section 5.1
@@
In addition to TOFU, we also evaluate the utility of the unlearned model on general task benchmarks. Performance on these tasks is also an important indicator of retain performance, reflecting whether general capabilities of LLMs are preserved after unlearning. As shown in Table 2, $\delta$-Unlearning achieves competitive performance on most metrics when compared to direct fine-tuning baselines, and closes the performance gap between before and after unlearning. $\delta$-Unlearning consistently outperforms direct fine-tuning with KL minimization, and bring improvement on 3 out of 4 tasks with gradient difference and data relabeling.
@@
Overall, our experiments demonstrate that $\delta$-Unlearning is a strong alternative to direct fine-tuning, with matching or even superior performance in terms of both forget and retain performance. $\delta$-Unlearning is also effective across different unlearning algorithms, showing the versatility of our approach.
5 Analysis
Figure 2: Unlearning trajectory of Gradient Ascent using direct fine-tuning (left), $\delta$-Unlearning (middle), and the tradeoff curve between forget and retain performance (right) over the course of unlearning. For training trajectories we report ROUGE score on all four TOFU datasets. For the tradeoff curve we report Forget Set ROUGE versus Non-forget Set ROUGE score.
5.1 Unlearning Trajectory
To better understand how forget and retain performance change over the course of unlearning, we study the training trajectory of both direct fine-tuning and $\delta$-Unlearning. As shown in Figure 2, when targeting the same ROUGE score on the forget set, $\delta$-Unlearning exhibits a steeper decline on the forget set initially compared to direct fine-tuning, which is also coupled with a steeper decline on the Retain Set. As unlearning progresses, direct fine-tuning starts to lose performance on the Real Author set that the model should not forget, while $\delta$-Unlearning still maintains relatively stable performance.
When comparing the training trajectory on different TOFU datasets, we can clearly observe the varying degrees of knowledge entanglement with the Forget Set. Being the most similar to the Forget Set, Retain Set performance starts to degrade at early stages, followed by the Real Author set. Performance on the World Fact, which is the least relevant to the Forget Set, only declines slightly towards the end of unlearning. This highlights the importance of finding a good balance between forget and retain performance for an unlearning method. We also study this trade-off from a more direct perspective by plotting the curve forget set ROUGE score versus non-forget set ROUGE score in Figure 2 (right). A desired model should lie at the upper right corner, which represents strong forget and retain performance. While Direct fine-tuning maintains more performance on retain data initially, $\delta$-Unlearning achieves a better balance at higher unlearning levels as direct fine-tuning starts to lose more performance on non-forget sets.
5.2 Effect of Offset Strength
As we mentioned in Section 3.2, we can adjust the value of $\alpha$ to control the strength of the logit offset being added to the larger model’s logits. We experiment with using difference $\alpha$ values during inference and study its effect on forget and retain performance. As shown in Figure 3. A low offset strength makes the effect of logit offset negligible, and the prediction of the ensemble is essentially dominated by the larger model $M$ without unlearning. As we gradually increase the offset strength, the unlearning effect becomes more prominent and forget set ROUGE score decreases significantly. Similar to what we observe in Figure 2, Retain Set performance largely follows the trajectory of Forget Set, while Real Author and World Fact performance are less influenced by the increase of unlearning offset strength. When we surpass the level of offset strength used in training, we observe continued performance degradation on all four datasets. The ROUGE score on forget set drops below 10 when $\alpha$ increases to 5, a score much lower than the retraining baseline (which has a ROUGE score of 38.9). However, at this offset strength the model becomes unusable, indicated by poor performance across all three non-forget set.
@@ 5.2 Effect of Offset Strength
$\alpha=0.5$ Hina Ameen primarily contributes to the genre of geology. Her extensive knowledge of …
$\alpha=1.0$ Hina Ameen works primarily in the genre of mythology. Her literature has a deep connection …
@@
We present an example from the Forget Set in Table 4 to provide a better understanding of model behavior at different offset strength levels. At $\alpha$=0.2, the model can perfectly reproduce the answer that is supposed to be forgotten, showing that unlearning is not taking effect at low offset strength. At $\alpha$=0.5, the model is still capable of recalling the correct answer, despite in slightly different phrasing. At $\alpha$=1, we obtain a fluent response with the sensitive information from the ground truth removed, demonstrating success of unlearning. If we further increase offset strength, the model starts to generate gibberish and eventually becomes unusable. In conclusion, using the same offset strength as in training leads to the best results overall.
5.3 Choice of Offset Model
Model TOFU Forget TOFU Retain
RL ($\downarrow$) P ($\downarrow$) TR ($\uparrow$) FQ ($\uparrow$) MU ($\uparrow$)
Retraining 37.2 11.6 57.7 1.0 58.8
7B Direct Fine-tuning 37.4 4.1 50.1 1.2e-4 37.8
7B + 3B Offset 37.3 37.0 46.7 1.5e-7 52.2
7B + 1.5B Offset 36.9 17.5 49.9 6.7e-6 49.1
7B + 0.5B Offset 37.0 14.7 43.8 4.9e-10 48.8
Table 5: Results on different offset model size.
To study how the choice of offset model affects unlearning performance, we run a series of experiments using Qwen family models as they offer a wide range of model sizes. Specifically, we apply offset unlearning with gradient ascent to a target Qwen2.5-7b-instruct model using 3B, 1.5B and 0.5B offset models. As shown in Table 5, all models can reach a level of unlearning similar to retraining as measured by ROUGE score of the generated responses, and using a stronger offset model generally leads to better performance on retain sets. Distributional metrics such as truth ratio and forget quality are more sensitive to the choice of offset model especially when the offset models are weak, as weak offset models tend to behave very differently from the target model during unlearning, and thus making the final logit ensemble deviate more from direct fine-tuning. Similar to what we observe in Table 3, using different offset models affects how the model balances performance across different types of knowledge during unlearning.
@@ 6 Conclusion
Acknowledgment
We appreciate the reviewers for their insightful comments and suggestions. James Y. Huang was supported by a gift fund from the USC Center on Secure & Trusted ML. Fei Wang was supported by an Amazon ML PhD Fellowship. Muhao Chen was supported by the DARPA FoundSci Grant HR00112490370, the NSF of the United States Grant ITE 2333736.